### v2

# P2 Law: Scaling Law for Post-Training After Model Pruning

Abstract
Pruning has become a widely adopted technique for reducing the hardware requirements of large language models (LLMs). To recover model performance after pruning, post-training is commonly employed to mitigate the resulting performance degradation. While post-training benefits from larger datasets, once the dataset size is already substantial, increasing the training data provides only limited performance gains. To balance post-training cost and model performance, it is necessary to explore the optimal amount of post-training data. Through extensive experiments on the Llama-3 and Qwen-2.5 series models, pruned using various common pruning methods, we uncover the scaling Law for Post-training after model Pruning, referred to as the P2 Law. This law identifies four key factors for predicting the pruned model’s post-training loss: the model size before pruning, the number of post-training tokens, the pruning rate, and the model’s loss before pruning. Moreover, P2 Law can generalize to larger dataset sizes, larger model sizes, and higher pruning rates, offering valuable insights for the post-training of pruned LLMs.
P2 Law: Scaling Law for Post-Training After Model Pruning
Xiaodong Chen1,2, Yuxuan Hu1,2, Xiaokang Zhang1,2, Yanling Wang3   Cuiping Li1,2Hong Chen1,2Jing Zhang1,2thanks: Corresponding author. 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3 Zhongguancun Laboratory, China {chenxiaodong,huyuxuan1999,zhang-jing,zhang2718,licuiping,chong}@ruc.edu.cn  wangyl@zgclab.edu.cn
1 Introduction
Large language models (LLMs) based on the Transformer architecture Vaswani et al. (2017) have been applied across diverse domains and tasks. However, as LLMs grow in size, their hardware demands increase substantially, limiting their practical deployment in real-world scenarios. To address this challenge, researchers have focused on developing compact models through model pruning techniques Han et al. (2016) that maintain high performance while reducing hardware requirements.
Figure 1: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15%. $C$ denotes the computational cost, which is calculated by $C=6ND$ Kaplan et al. (2020), where $N$ denotes the model size after pruning, and $D$ denotes the number of post-training tokens.
Model pruning can be broadly categorized into unstructured pruning  Frantar and Alistarh (2023); Zhang et al. (2024); Sun et al. (2024) and structured pruning Chen et al. (2024); Hu et al. (2024); Liu et al. (2024); Muralidharan et al. (2024); Ma et al. (2023); Ashkboos et al. (2024); Men et al. (2024). Unstructured pruning removes individual elements from weight matrices, producing sparse matrices while preserving satisfactory model performance. However, the introduced structural irregularities make this approach hardware-unfriendly and hinder its ability to accelerate computation. To mitigate this problem, semi-structured pruning, a variant of unstructured pruning, leverages specific hardware support Mishra et al. (2021) to achieve acceleration but may result in greater performance degradation compared to unstructured pruning. In contrast, structured pruning removes entire components, such as attention heads or layers, effectively reducing the model size but often with a higher performance drop compared to other pruning methods.
To effectively leverage hardware-friendly models pruned using semi-structured or structured pruning methods, post-training Ashkboos et al. (2024); Chen et al. (2024); Yang et al. (2024); Ma et al. (2023); Kim et al. (2024) serves as an essential step after model pruning to mitigate the performance degradation. For example, LLM-Pruner Ma et al. (2023) utilizes 50,000 instruction data samples for fine-tuning, whereas Shortened Llama Kim et al. (2024) uses 627B tokens of pre-training data for continual pre-training of the pruned LLMs. In general, compared to fine-tuning with a small dataset, continual pre-training with a large dataset is a more effective way to fully recover performance, but it demands substantial hardware resources. Given the significant hardware demands, a question is raised: is it truly necessary to use a vast amount of data for performance recovery? LLM-Streamline Chen et al. (2024) answers the question by demonstrating that using large amounts of data for post-training only slightly improves performance compared to using a suitably sized amount. Hence, this raises another question: whether a scaling law can be established to predict the optimal amount of post-training data required after model pruning for resource efficiency?
To address the problem, we conduct pilot experiments on the Llama-3 Dubey et al. (2024) and Qwen-2.5 series models Team (2024), applying both typical structured and semi-structured pruning methods. In specific, we observe several trends in the post-training loss curves, allowing us to identify the necessary conditions that the scaling Law for Post-training after model Pruning (P2 Law) must satisfy. Building on the Chinchilla scaling law Hoffmann et al. (2022) proposed for pre-training and the identified conditions, we define multiple parameterizations of our P2 Law and select the most suitable parameterization. To assess the fit of different parameterizations to P2 Law, we introduce a new metric named Average Slope Difference (ADS). As scaling laws are used to find the suitable training data size by balancing cost and performance, focusing on the slope of the predicted loss curve rather than the predicted loss values, the ADS metric is designed to measure the slope discrepancy between predicted and actual loss curves. Finally, P2 Law is parameterized as,
$\begin{split}\mathcal{L}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (1)
where $N_{C}$, $D_{C}$, $E$, $\alpha$, $\beta$, $\gamma$, $\delta$ are constants, $N_{0}$ denotes the model size before pruning, $D$ denotes the number of post-training tokens, $\rho$ denotes the pruning rate, $\mathcal{L}_{0}$ denotes the model’s loss before pruning, and $\mathcal{L}$ denotes the pruned model’s post-training loss.
In this paper, we conduct a series of experiments to validate the P2 Law. Taking Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15% as an example, Figure 1 illustrates that P2 Law accurately predicts the actual post-training losses of the pruned model checkpoints, where $C$ represents the computational cost calculated as $C=6ND$ Kaplan et al. (2020), $N$ is the model size after pruning, and $D$ is the number of post-training tokens. Utilizing the post-training loss curves derived by P2 Law, we can accurately predict that the computational cost required for the post-training loss of Llama-3.2-1B to start decreasing gently is approximately $10^{4}$. This predicted size of post-training data provides a good balance between cost and performance. Furthermore, we evaluate the generalization ability of the P2 Law, demonstrating that P2 Law can effectively generalizes to larger dataset sizes, larger model sizes, and higher pruning rates.
Overall, this work makes the following contributions:
  • We conduct extensive studies to uncover the P2 Law, the first scaling law for post-training after pruning, helping balance post-training cost and pruned LLM performance.
  • We propose ADS, an effective metric for the evaluation of parameterizations of scaling laws for the post-training of pruned LLMs.
  • We demonstrate that the P2 Law generalizes effectively to larger dataset sizes, larger models, and higher pruning rates, offering valuable insights for optimizing pruned LLMs across diverse settings.
2 Preliminary
In this section, we present the preliminary of this work, including various pruning methods and the post-training method.
2.1 Pruning
We utilize three common pruning methods to prune LLMs, including two structured pruning methods (depth pruning Chen et al. (2024); Song et al. (2024); Gromov et al. (2024) and width pruning Ashkboos et al. (2024); Hu et al. (2024); Liu et al. (2024)) and a hardware-friendly variant of unstructured pruning method known as 2:4 semi-structured pruning Sun et al. (2024); Frantar and Alistarh (2023); Zhang et al. (2024).
Depth Pruning. Depth pruning is a structured pruning method that removes entire Transformer layers from LLMs. Specifically, depth pruning involves estimating the importance of each Transformer layers in LLMs and then removing those layers with the lowest importance.
Width Pruning. Width pruning is another structured pruning method that reduces the number of embedding channels in LLMs. This method involves measuring the importance of embedding channels and pruning the least important ones.
2:4 Semi-Structured Pruning. Unstructured pruning removes individual unimportant elements from the weight matrices, producing sparse matrices. 2:4 semi-structured pruning is a variant of unstructured pruning, with a sparse pattern of 2:4. In this pattern, every four elements in the weight matrices are grouped together, with two of the elements in each group set to zero. This semi-structured sparsity can be efficiently accelerated by hardware. We utilize SparseGPT Frantar and Alistarh (2023), a well-known 2:4 semi-structured pruning method, to prune LLMs.
For more details about the pruning methods used in this paper, please refer to the Appendix B.
2.2 Post-Training
After the pruning, we conduct post-training on the pruned LLMs to mitigate the performance decline. For LLMs pruned using depth or width pruning, we train all parameters of the pruned LLMs. For sparse LLMs derived from 2:4 semi-structured pruning, inspired by SP-LoRA Anonymous (2024), we combine the updated weight from each training iterate with the sparse mask during the post-training process to ensure the model’s sparsity, further post-training details about the 2:4 semi-structured pruning is provided in Appendix B.3.
3 Experiments for Finding Necessary Conditions Satisfied by P2 Law
In this section, we conduct experiments on six LLMs from the Llama-3 and Qwen-2.5 series, covering various model sizes and using depth pruning, width pruning, and 2:4 semi-structured pruning.
First, we detail the pruning settings and post-training settings in Section 3.1. Next, we describe multiple trends observed in the post-training loss curves in Section 3.2. Finally, in Section 3.3, we identify several necessary conditions that the P2 Law must satisfy based on the observed trends.
(a) Post-training loss curves of Llama-3.2-1B pruned by depth pruning with different pruning rates.
(b) Post-training loss curves of Llama-3.2-3B pruned by depth pruning with different pruning rates.
(c) Post-training loss curves of Llama-3.1-8B pruned by depth pruning with different pruning rates.
Figure 2: Post-training loss curves of Llama-3 series models pruned by depth pruning with different pruning rates.
Figure 3: Post-training loss curves of Llama-3 series models pruned by 2:4 semi-structured pruning.
Figure 4: Normalized relative post-training loss curves of Llama-3.1-8B pruned by depth pruning.
Depth pruning Width pruning
Llama-3.2-1B 15%,25%,30% 15%,25%,35%
Llama-3.2-3B 16%,25%,34% 15%,25%,35%
Llama-3.1-8B 16%,24%,33% 15%,25%,35%
Qwen-2.5-0.5B 15%,21%,27% 15%,25%,35%
Qwen-2.5-1.5B 15%,24%,33% 15%,25%,35%
Qwen-2.5-3B 17%,25%,32% 15%,25%,35%
Table 1: Pruning rates used for depth pruning and width pruning on different LLMs.
3.1 Settings
We conduct experiments on six LLMs from the Llama-3 and Qwen-2.5 series, including Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, Qwen-2.5-0.5B, Qwen-2.5-1.5B and Qwen-2.5-3B.
Pruning. The pruning rates used for depth pruning and width pruning are shown in Table 1. The pruning processes have been introduced in Appendix B. We randomly select 1,024 samples from the pre-training dataset SlimPajama for pruning.
Post-Training. For Llama-3.2-3B, Qwen-2.5-3B and Llama-3.1-8B, we randomly select 1B tokens from SlimPajama for post-training. For Llama-3.2-1B, Qwen-2.5-0.5B, Qwen-2.5-1.5B, we randomly select 0.5B tokens from SlimPajama for post-training. During the post-training process, we set the learning rate to 2e-5 and the batch size to 256k tokens. All post-training processes are conducted on 4 Nvidia A800-80G GPUs and 4 Nvidia A6000-48G GPUs. The entire training process takes a total of 500 hours.
3.2 Trends of the Post-Training Loss Curves
To better explore the trends of the post-training loss curves, we define:
Definition 1
Relative post-training loss $\Delta\mathcal{L}$. The relative post-training loss is the difference between the pruned model’s post-training loss $\mathcal{L}$ and the model’s loss $\mathcal{L}_{0}$ before pruning.
$\begin{split}\Delta\mathcal{L}=\mathcal{L}-\mathcal{L}_{0}\end{split}$ (2)
Definition 2
Normalized relative post-training loss $\Delta\mathcal{L}_{norm}$. The normalized relative post-training loss is defined as the ratio of the relative post-training loss $\Delta\mathcal{L}$ to a power-law function of the pruning rate $\rho$.
$\begin{split}\Delta\mathcal{L}_{norm}=\frac{\Delta\mathcal{L}}{(\frac{1}{\rho})^{\gamma}}\end{split}$ (3)
where $\gamma$ is a constant.
In Figures 2 and 4, we present the post-training loss curves for the Llama-3 series models pruned by depth pruning and 2:4 semi-structured pruning. Additional post-training loss curves (exhibiting similar trends) are shown in Figures 101112, and 14 in the Appendix C. By analyzing the post-training loss curves, we observe the following trends:
  • Trend 1: Smaller LLMs exhibit faster decreases in post-training loss. For instance, as shown in Figure 4, with 2:4 semi-structured pruning, the post training loss curve of Llama-3.1-8B is much flatter compared to those of Llama-3.2-3B and Llama-3.2-1B. The same trend is observed under both depth pruning and width pruning, as depicted in Figure 2 and Figure 10. This suggests that smaller LLMs exhibit faster decreases in post-training loss.
  • Trend 2: Relative post-training loss $\Delta\mathcal{L}$ follows a power-law relationship with the pruning rate $\rho$. As shown in Figure 4, with depth pruning, the normalized relative post-training loss curves of Llama-3.1-8B at various pruning rates nearly overlap. This can be formally expressed as:
    $\begin{split}\frac{\Delta\mathcal{L}^{(0.33)}}{(\frac{1}{0.33})^{\gamma}}\approx\frac{\Delta\mathcal{L}^{(0.24)}}{(\frac{1}{0.24})^{\gamma}}\approx\frac{\Delta\mathcal{L}^{(0.16)}}{(\frac{1}{0.16})^{\gamma}}\end{split}$ (4)
    where $\Delta\mathcal{L}^{(0.33)}$, $\Delta\mathcal{L}^{(0.24)}$, and $\Delta\mathcal{L}^{(0.16)}$ represent the relative post-training loss of Llama-3.1-8B pruned by depth pruning with pruning rates $\rho$ of 0.33, 0.24, and 0.16, respectively. This demonstrates that the pruning rate and the relative post-training loss are governed by a power-law relationship.
3.3 Necessary Conditions Satisfied by P2 Law
Based on the aforementioned trends, we identify three fundamental conditions for the P2 Law:
  • Condition 1. The post-training loss $\mathcal{L}$ decreases as the number of post-training tokens $D$ increases:
    $\begin{split}\frac{\partial\mathcal{L}}{\partial D}<0\end{split}$ (5)
  • Condition 2. As derived from Trend 1 in Section 3.2, under similar pruning rates, the post-training loss curves of smaller LLMs decrease faster as the number of post-training tokens increases:
    $\begin{split}\frac{\partial}{\partial N_{0}}(\frac{\partial\mathcal{L}}{\partial D})=\frac{\partial^{2}\mathcal{L}}{\partial N_{0}\partial D}>0\end{split}$ (6)
    where $N_{0}$ is the model size before pruning.
  • Condition 3. From Eq.4 in Trend 2, the relative post-training loss $\Delta\mathcal{L}$ follows a power-law relationship with the pruning rate $\rho$.:
    $\begin{split}\Delta\mathcal{L}\propto(\frac{1}{\rho})^{\gamma}\end{split}$ (7)
An ideal P2 Law should satisfy aforementioned three conditions. Additionally, the P2 Law should also satisfy the condition that when the pruning rate $\rho$ is 0, the relative post-training loss $\Delta\mathcal{L}$ is 0, which is a necessary condition for Condition 3.
4 P2 Law
In this section, we aim to parameterize the P2 Law according to the above three necessary conditions. In Section 4.1, we introduce the metric for assess the quality of different candidate parametrizations. Next, based on the Chinchilla scaling law, we define multiple parameterizations for our P2 Law and select the most suitable one in Section 4.2. Finally, in Section 4.3, we demonstrate the generalization ability of the P2 Law from three perspectives: dataset size, model size, and pruning rate.
4.1 Metric for Accessing Law Fitting
Following prior work Que et al. (2024), we utilize both $R^{2}$ Fisher (1922) and Huber loss Huber (1992) to evaluate different parameterizations of scaling law. The $R^{2}$ value, or coefficient of determination, measures the fitting quality, with values closer to 1 indicating a better fit. Huber loss, a robust loss function, blends the characteristics of mean squared error and mean absolute error, making it less sensitive to outliers. The Huber loss is a positive number, and a lower Huber loss suggests a better fit.
Scaling laws are often used to determine the optimal amount of training data by balancing computational cost and model performance. For instance, as shown in Figure 5, there is one actual loss curve and two predicted loss curves. Traditional metrics like $R^{2}$ and Huber loss indicate that predicted curve 1 better matches the actual curve. However, when predicted curve 1 is used to predict the point where the actual curve starts to flatten, it results in a large error. On the other hand, while predicted curve 2 deviates more from the actual curve in terms of absolute values, its slope is consistently closer to the actual curve. This makes its prediction of the flattening point much more accurate. To address this issue, we propose a new metric called Average Slope Difference (ASD), which measures the difference between the slope of the loss curve predicted by the scaling law and the slope of the actual loss curve. ASD is formally defined as:
$\text{ASD}=\frac{1}{N}\sum_{i=2}^{N}\left|(y_{i}-y_{i-1})-(\hat{y_{i}}-\hat{y_{i-1}})\right|$ (8)
where ${y_{i}}$ represents the loss of $N$ points uniformly sampled from the actual loss curve as the number of post-training tokens increases, and $\hat{y_{i}}$ represents the corresponding loss values on the curve predicted by the scaling law. Since the early parts of the loss curve during post-training do not represent true convergence, we only sample points from the latter half of the training process. A smaller ASD value indicates that the predicted loss curve’s slope more closely matches the slope of the actual loss curve.
LLM Parameterizations Depth pruning Width pruning 2:4 semi-structured pruning
$R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD
Llama-3 series $\mathcal{L}_{1}$ 0.9717 0.000016 0.000619 -1.2985 0.000177 0.000592 0.8126 0.000056 0.001466
$\mathcal{L}_{2}$ 0.9300 0.000045 0.001150 -2.5578 0.000450 0.001419 0.7797 0.000079 0.002294
$\mathcal{L}_{3}$ 0.7737 0.000118 0.000827 -4.5905 0.000776 0.001754 -0.2555 0.000493 0.002054
Qwen-2.5 series $\mathcal{L}_{1}$ 0.9781 0.000011 0.000524 0.9891 0.000010 0.000648 0.9995 0.000000 0.000191
$\mathcal{L}_{2}$ 0.9423 0.000031 0.000879 0.9803 0.000027 0.000712 0.9867 0.000010 0.000753
$\mathcal{L}_{3}$ 0.8855 0.000075 0.001270 0.9824 0.000024 0.000733 0.9930 0.000005 0.000491
Table 2: Evaluation of three parameterizations for P2 Law fitting.
Figure 5: A example showcasing the advantages of ADS. The ADS of predicted loss curve 2 is lower because its slope is closer to that of the actual loss curve.
(a) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.2-1B pruned by depth pruning.
(b) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.2-3B pruned by depth pruning.
(c) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.1-8B pruned by depth pruning.
Figure 6: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series model pruned by depth pruning method.
4.2 Derivation of P2 Law
Previous efforts have explored scaling laws for pre-training of LLMs, with Chinchilla scaling Hoffmann et al. (2022) being a superior work, and we choose it as the foundational parameterization for our P2 Law. The Chinchilla scaling law describes the relationship between model performance and key factors such as model size, the number of pre-training tokens, and the computational resources used during the pre-training process. It is formally defined as follows:
$\begin{split}\mathcal{L}(N,D)=\frac{N_{C}}{N^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E\end{split}$ (9)
where $N_{C}$, $D_{C}$, $E$, $\alpha$, and $\beta$ are constants, $N$ represents the model size, $D$ denotes the number of pre-training tokens and $\mathcal{L}$ represents the model’s loss. Compared to the OpenAI scaling law Kaplan et al. (2020), the Chinchilla scaling law demonstrates superior performance (detailed in Appendix D). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law. Combining the pruning rate $\rho$ and the model’s loss $\mathcal{L}_{0}$ before pruning, we define the following three candidate parameterizations:
$\begin{split}\mathcal{L}_{1}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (10)
$\begin{split}\mathcal{L}_{2}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (11)
$\begin{split}\mathcal{L}_{3}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (12)
where $N_{C}$, $D_{C}$, $E$, $\alpha$, $\beta$, $\gamma$ and $\delta$ are constants, $N_{0}$ denotes the model size before pruning, $D$ denotes the number of post-training tokens and $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ denote the pruned model’s post-training loss. Additionally, since there is no pruning rate in the 2:4 semi-structured pruning, P2 Law for the 2:4 semi-structured pruning does not need to satisfy Condition 3, and we adjust the parameterizations to:
$\begin{split}\mathcal{L}_{1}(N_{0},D)=(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (13)
$\begin{split}\mathcal{L}_{2}(N_{0},D)=(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (14)
$\begin{split}\mathcal{L}_{3}(N_{0},D)=(\frac{1}{N_{0}})^{\delta}(\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (15)
LLM Generalization Depth pruning Width pruning 2:4 semi-structured pruning
$R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD
Llama-3 series Dataset size 0.9725 0.000016 0.001001 0.9270 0.000019 0.000674 0.8244 0.000063 0.002561
Model size -0.5441 0.000745 0.001321 - - - - - -
Pruning rate 0.9676 0.000059 0.000879 0.9707 0.000056 0.001123 - - -
Qwen-2.5 series Dataset size 0.9780 0.000012 0.001026 0.9896 0.000010 0.001116 0.9940 0.000000 0.000299
Model size -0.8786 0.000763 0.001573 0.8627 0.000137 0.001772 - - -
Pruning rate 0.9660 0.000043 0.000920 0.9704 0.000095 0.001003 - - -
Table 3: Evaluation of generalization results from the perspectives of dataset size, model size, and pruning rate.
Figure 7: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by width pruning).
Figure 8: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by width pruning).
Figure 9: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by width pruning).
We utilize all checkpoints to fit the three candidate parameterizations through Levenberg-Marquardt method Moré (2006), and the specific parameter values (i.e., the values of $N_{c}$, $D_{c}$, $E$, $\alpha$, $\beta$, $\gamma$, and $\delta$) for the fitted $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ are provided in Table 5 in Appendix E. As shown in Table 2, $\mathcal{L}_{1}$ significantly outperforms $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ in terms of the $R^{2}$, Huber loss, and ASD metrics. Additionally, as shown in Table 6 in Appendix E, after our calculation and verification, $\mathcal{L}_{2}$ and some fitted $\mathcal{L}_{3}$ fails to satisfy Condition 2. In contrast, all of the fitted $\mathcal{L}_{1}$ satisfy all three conditions. Based on the experimental results, we select $\mathcal{L}_{1}$ as the parameterization for our P2 Law.
In Figure 6, we show the loss curves $\mathcal{L}_{1}$ derived by P2 Law alongside the actual checkpoints of the Llama-3 series models pruned by depth pruning, where the computational cost ($C$) is approximated using the empirical formula $C=6ND$ Kaplan et al. (2020), and $N$ denotes the model size after pruning. Additional loss curve derived by P2 Law are shown in Figure 151617 and 18 in Appendix F. As shown in these figures, the loss curve derived by P2 Law accurately aligns with all actual checkpoints, under all the three pruning methods, except for the width pruning on Llama-3.1-8B. As shown in Figure 2(c) and 10(c), for Llama-3.1-8B, we observe that depth pruning outperforms width pruning at similar pruning rates, which contrasts with the observations in other cases. This suggests that width pruning on Llama-3.1-8B may lead to anomalous performance, making our law unsuitable for this special scenario. We elaborate on this anomalous performance of width pruning on Llama-3.1-8B further in Appendix G.
4.3 Generalization of P2 Law
In this section, we explore the generalization ability of P2 Law from three perspectives: dataset size, model size and pruning rate.
4.3.1 Settings
We begin by outlining the settings of generalization experiments as follows:
Dataset Size. The fitting setting follows the same setting as described in Section 4.2, with the only difference being that the first 80% of the checkpoints recorded during each training process are used to fit the P2 law, and the remaining 20% for validation.
Model size. We fit the P2 Law using checkpoints from smaller LLMs, including Llama-3.2-1B, Llama-3.2-3B, Qwen-2.5-0.5B, and Qwen-2.5-1.5B. We then validate the fitted P2 Law on checkpoints from larger LLMs, including Llama-3.1-8B and Qwen-2.5-3B. Taking the Qwen-2.5 series models as an example, we fit the P2 Law using all checkpoints from Qwen-2.5-0.5B and Qwen-2.5-1.5B, and subsequently validate it with the actual checkpoints of Qwen-2.5-3B across three pruning rates. Due to the limited number of available actual loss curves for 2:4 semi-structured pruning, we did not conduct experiments for this method.
Pruning Rate. We fit the P2 Law using checkpoints from lower pruning rates and then validate it with checkpoints from higher pruning rate. Taking width pruning of the Qwen-2.5 series models as an example, we fit the P2 Law using checkpoints from these models at lower pruning rates (0.15 and 0.25) and then validate it with the actual checkpoints at a higher pruning rate of 0.35. Since there is no pruning rate in the 2:4 semi-structured pruning, we only explore the generalization ability on pruning rates under depth pruning and width pruning.
Due to the anomaly of the width pruning on Llama-3.1-8B (detailed in the last paragraph of Section 4.2), we do not consider the generalization experiments on this model with width pruning.
4.3.2 Dataset Size Generalization
The evaluation results are shown in Table 3, and the loss curves of Qwen-2.5-3B (pruned by width pruning) derived by P2 Law are illustrated in Figure 9. Additional loss curves derived by P2 Law are provided in Figures 202122, and 23 in Appendix H. The results in Table 3 show that the loss curves derived by P2 Law accurately matches the validation checkpoints, indicating that the P2 Law generalizes well to larger dataset sizes.
4.3.3 Model Size Generalization
The evaluation results are presented in Table 3, and the loss curves of Qwen-2.5-3B predicted by P2 Law (pruned by width pruning) are visualized in Figure 9. Additional loss curves predicted by P2 Law are shown in Figure 25 in Appendix H. As shown in Table 3, the P2 Law fitted on smaller LLMs performs poorly in $R^{2}$ and Huber loss when applied to larger models, indicating challenges in generalizing to larger, unseen models. However, the low ASD suggests it still captures the slope of the actual loss curve. This trend is also seen in Figure 25, where despite a gap between predicted and actual loss curves, the predicted and actual loss curves align in their downward trend after training stabilizes. This suggests P2 Law fitted from smaller LLMs can still predict the optimal computation cost point for larger LLMs, confirming its generalization feasibility.
4.3.4 Pruning Rate Generalization
We present the generalization evaluations in Table 3 and illustrate the loss curves of Qwen-2.5 series models predicted by P2 Law (pruned by width pruning) in Figure 9. Additional loss curves predicted by P2 Law are provided in Figure 26 and 27 in Appendix H. As shown in the Figure 9 and Table 3, the values of different metrics indicate that the actual loss curves closely align with the predicted loss curves, suggesting that the P2 Law generalizes well to higher pruning rates.
5 Related Work
5.1 Model Pruning
Model pruning can be categorized into unstructured pruning and structured pruning.
Unstructured Pruning. Unstructured pruning methods Frantar and Alistarh (2023); Zhang et al. (2024); Sun et al. (2024) compress LLMs by removing individual unimportant elements from the weight matrices, producing sparse ones. However, it is often hardware-inefficient and only speeds up LLMs when a specific sparsity pattern, such as 2:4 sparsity Mishra et al. (2021), is applied. The approach which employ the 2:4 sparsity is known as semi-structured pruning.
Structured Pruning. Structured pruning methods for LLMs can be divided into two categories: depth pruning Chen et al. (2024); Song et al. (2024); Gromov et al. (2024); Men et al. (2024), which aims to reduce the number of layers in the LLMs, and width pruning Ashkboos et al. (2024); Hu et al. (2024); Liu et al. (2024); Ma et al. (2023), which aims to reduce the embedding channels, the number of attention heads, or the intermediate size of the FFN.
5.2 Scaling Law
The OpenAI scaling law Kaplan et al. (2020) and the Chinchilla scaling law Hoffmann et al. (2022) are the most popular scaling laws in the pre-training of LLMs, both of which establishe a power-law relationship between model performance, model size, the number of pre-training tokens, and the computational resources used during pre-training.
We are the first to investigate the scaling law for the post-training after model pruning, and we propose the P2 Law as a scaling law for this process. Compared to the OpenAI scaling law, the Chinchilla scaling law demonstrates superior performance (detailed in Appendix D). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law.
6 Conclusion
In this paper, we conduct post-training experiments on models from the Llama-3 and Qwen-2.5 series, covering various sizes and employing both typical structured and semi-structured pruning methods. Through extensive experiments, we identify the P2 Law — the first scaling law for post-training after model pruning. Further experiments validate the effectiveness of the P2 Law and demonstrate its generalization to larger dataset sizes, larger model sizes, and higher pruning rates, offering valuable insights for resource allocation in the post-training of pruned LLMs.
Limitation
The P2 Law presented in this paper, derived from experiments on the Llama-3 and Qwen-2.5 series models, may not be directly applicable to other architectures, such as the currently popular Mixture of Experts (MoE) architecture. Therefore, we plan to conduct more experiments across a broader range of model architectures to enhance the generalizability of our findings.
Appendix A License
Our research is grounded in the SlimPajama training dataset, which is distributed under the Apache 2.0 license. This license allows for the free use, modification, reproduction, and distribution of the software, both for personal and commercial purposes. Consistent with open science practices, we will make our training data publicly available upon acceptance of this work. The data will be released under the CC BY-SA 4.0 license, which enables reuse and redistribution, provided that derivative works adhere to the same licensing terms
Appendix B Details of Pruning Methods
B.1 Depth Pruning
Following the existing depth pruning methods Men et al. (2024); Chen et al. (2024); Yang et al. (2024), we estimate the layer importance using cosine similarity and prune layers with lower importance. Specifically, we randomly select $N$ samples from the pre-training data. We then record the hidden states generated by the LLMs for these samples and compute the cosine similarity between the input and output hidden states of each layer. Assuming that the input hidden states of layer $i$ are represented by $\boldsymbol{x}^{(i)}$, the importance score(IS) of layer $i$ is computed as:
$\text{IS}^{\text{layer},i}=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{1}{L}\sum_{k=1}^{L}\frac{\boldsymbol{x}^{(i)}_{j,k}\cdot\boldsymbol{x}^{(i+1)}_{j,k}}{\|\boldsymbol{x}^{(i)}_{j,k}\|\cdot\|\boldsymbol{x}^{(i+1)}_{j,k}\|}\right)$ (16)
where $\boldsymbol{x}^{(i)}_{j}$, $\boldsymbol{x}^{(i+1)}_{j}\in\mathbb{R}^{d\times L}$ denotes the input and output hidden states of the $j$-th sample respectively, $L$ denotes the sequence length and $d$ denotes the hidden size. Given the number of pruned layers $n$ determined by the target sparsity, we remove the $n$ layers corresponding to the top-$n$ highest cosine similarities for pruning.
B.2 Width Pruning
Following the approaches of Wanda Sun et al. (2024) and MINITRON Muralidharan et al. (2024), we utilize activation-based metrics for width pruning. Specifically, we randomly select $N$ samples from the pre-training data and assess the importance of embedding channels by analyzing the activations generated by the LayerNorm layers. We then prune the least important channels based on this analysis. The formula for calculating the importance score (IS) of embedding channels (emb) is as follows:
$\text{IS}^{\text{emb},i}=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{1}{L}\sum_{k=1}^{L}\left|LN(\boldsymbol{x}_{j,k,i}^{LN})\right|\right)$ (17)
where $\boldsymbol{x}_{j,k,i}^{LN}$ denotes the input of the $i$-th channel of the $k$-th token in the $j$-th sample at the LayerNorm layer, $L$ denotes the sequence length, and $LN$ denotes the Layer Normalization operaten. Given a specific sparsity, we calculate the number of embedding channels that need to be pruned, and then remove the channels with the lowest importance.
B.3 2:4 Semi-Structured Pruning
Unstructured pruning removes individual unimportant elements from the weight matrices, producing sparse matrices. When the sparsity structure follows a specific pattern, such as 2:4 sparsity Mishra et al. (2021), the model can be efficiently accelerated. This approach is known as semi-structured pruning. Let $W$ represent the weight matrix of a linear layer of an LLM, $x$ represent the input of the linear layer. The object of semi-structured pruning is to learn a sparsity mask $M$ and an updated weight $\Delta W$ so that the dense matrix $W$ is transformed into a sparse matrix $\tilde{W}$:
$min\ \ \|Wx-\tilde{W}x\|$
$s.t.\quad\tilde{W}=M\cdot(W+\Delta W)$ (18)
where $W\in\mathbb{R}^{d_{out}\times d_{in}}$, $M\in\{0,1\}^{d_{out}\times d_{in}}$, $\Delta W\in\mathbb{R}^{d_{out}\times d_{in}}$ and $x\in\mathbb{R}^{d_{in}}$.
We randomly select 1,024 data samples from the pre-training dataset SlimPajama for pruning. and use SparseGPT Frantar and Alistarh (2023) to optimize the aforementioned objectives.
In the post-training process, We train this 2:4 sparse model pruned by SparseGPT. Inspired by SP-LoRA Anonymous (2024), during the post-training process, we combine the updated weight $\Delta\tilde{W}^{t}$ from each training iterate $t$ with the mask $M$ to obtain the weight after update $\tilde{W}^{t}$, ensuring the model’s sparsity:
$\tilde{W}^{t}=\tilde{W}^{t-1}+M\cdot\Delta\tilde{W}^{t}$ (19)
Appendix C Additional Actual Loss Curves
The additional post-training loss curves for models pruned by width pruning or for the Qwen-2.5 series models are provided in Figures 1011, 12 and 14.
Appendix D Comparison with OpenAI Scaling Law
Kaplan Kaplan et al. (2020) propose OpenAI scaling law as follows:
$\begin{split}\mathcal{L}(N,D)=(\frac{N_{C}}{N^{\alpha}}+\frac{D_{C}}{D})^{\beta}\end{split}$ (20)
where $N_{C}$, $D_{c}$, $\alpha$ and $\beta$ are constants, $N$ denotes the model size and $D$ denotes the number of pre-training tokens. We have also defined the following parameterizations based on the OpenAI scaling law:
$\begin{split}\mathcal{L}_{4}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D})^{\beta}\end{split}$ (21)
$\begin{split}\mathcal{L}_{5}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D})^{\beta}\end{split}$ (22)
where $N_{C}$, $D_{C}$, $\alpha$, $\beta$, $\gamma$, $\delta$ denotes constants, $N_{0}$ denotes the model size before pruning, $D$ denotes the number of post-training tokens, $\rho$ denotes pruning rate, $\mathcal{L}_{0}$ denotes the model’s loss before pruning and $\mathcal{L}_{4},\mathcal{L}_{5}$ denote pruned model’s post-training loss.
We utilize all the checkpoints to fit the two parameterizations described above, and the evaluation results are presented in Table 4. The results show that the performance of these two parameterizations is weaker than that of $\mathcal{L}_{1}$. Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law.
Appendix E Parameter Values of Fitted Parameterizations
We present the parameter values of the fitted $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ in the Table 5. In addition, we calculate whether $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ satisfy Condition 2, and the results are shown in the Table 6.
Appendix F Additional Loss Curves Derived by P2 Law
The additional loss curves derived by P2 Law are shown in the Figure 151617 and 18.
Appendix G Patterns of the Llama-3 Series Models in Terms of Width
As discussed in Section 4.2, we observe an anomalous phenomenon in Llama-3.1-8B under width pruning. To investigate this further, we analyze the behavior of the Llama-3 series models with respect to width. Using a random sample of 1024 data points from SlimPajama and applying Eq.17, we plot the importance score distributions of the embedding channels for the Llama-3 series models, as shown in Figure 19. For easier comparison, we normalize the Importance score, which is defined as follows:
$\text{IS}^{\text{emb},i}=\frac{\text{IS}^{\text{emb},i}-\rm min(\text{IS}^{\text{emb},1},\text{IS}^{\text{emb},2},...,\text{IS}^{\text{emb},N_{d}})}{\rm max(\text{IS}^{\text{emb},1},\text{IS}^{\text{emb},2},...,\text{IS}^{\text{emb},N_{d}})}$
where the $N_{d}$ denotes the number of embedding channels. Additionally, we remove the extremely high values that represent a very small proportion of the data. The figure shows that the importance scores of Llama-3.1-8B are more densely distributed compared to those of Llama-3.2-1B and Llama-3.2-3B. This denser distribution may hinder the ability to effectively distinguish less important channels in Llama-3.1-8B based on importance scores, which could potentially explain the observed anomalies in Llama-3.1-8B.
Appendix H Additional Generalization Loss Curves
We present the additional dataset size generalization predicted loss curves in the Figure 20212223 and 24, model size generalization predicted loss curves in the Figure 25 and pruning rate generalization predicted loss curves in the Figure 26 and 27.
(a) Post-training loss curves of Llama-3.2-1B pruned by width pruning with different pruning rates.
(b) Post-training loss curves of Llama-3.2-3B pruned by width pruning with different pruning rates.
(c) Post-training loss curves of Llama-3.1-8B pruned by width pruning with different pruning rates.
Figure 10: Post-training loss curves of Llama-3 series models pruned by width pruning with different pruning rates.
(a) Post-training loss curves of Qwen-2.5-0.5B pruned by depth pruning with different pruning rates.
(b) Post-training loss curves of Qwen-2.5-1.5B pruned by depth pruning with different pruning rates.
(c) Post-training loss curves of Qwen-2.5-3B pruned by depth pruning with different pruning rates.
Figure 11: Post-training loss curves of Qwen-2.5 series models pruned by depth pruning with different pruning rates.
(a) Post-training loss curves of Qwen-2.5-0.5B pruned by width pruning with different pruning rates.
(b) Post-training loss curves of Qwen-2.5-1.5B pruned by width pruning with different pruning rates.
(c) Post-training loss curves of Qwen-2.5-3B pruned by width pruning with different pruning rates.
Figure 12: Post-training loss curves of Qwen-2.5 series models pruned by width pruning with different pruning rates.
Figure 13: Post-training loss curves of Qwen-2.5 series models pruned by 2:4 semi-structured pruning.
Figure 14: Post-training loss curves of Llama-3 series models pruned by 2:4 semi-structured pruning.
LLM Parameterizations Depth pruning Width pruning 2:4 semi-structured pruning
$R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD $R^{2}$ Huber loss ASD
Llama-3 series $\mathcal{L}_{1}$ 0.9717 0.000016 0.000619 -1.2985 0.000177 0.000592 0.8126 0.000056 0.001466
$\mathcal{L}_{4}$ 0.9339 0.000035 0.001482 -1.3660 0.000203 0.000814 0.7157 0.000112 0.002117
$\mathcal{L}_{5}$ 0.6535 0.000198 0.001822 -1.7948 0.000345 0.000729 -0.3809 0.000638 0.003687
Qwen-2.5 series $\mathcal{L}_{1}$ 0.9781 0.000011 0.000524 0.9891 0.000010 0.000648 0.9995 0.000000 0.000191
$\mathcal{L}_{4}$ 0.7730 0.000192 0.004085 0.9838 0.000015 0.001126 0.9960 0.000002 0.000550
$\mathcal{L}_{5}$ 0.8283 0.000134 0.002648 0.9694 0.000040 0.001007 0.8360 0.000118 0.003925
Table 4: Comparison of law fitting results between OpenAI scaling law and Chinchilla scaling law.
LLM Parameterizations Depth pruning Width pruning 2:4 semi-structured pruning
$N_{c}$ $D_{c}$ $E$ $\alpha$ $\beta$ $\gamma$ $\delta$ $N_{c}$ $D_{c}$ $E$ $\alpha$ $\beta$ $\gamma$ $\delta$ $N_{c}$ $D_{c}$ $E$ $\alpha$ $\beta$ $\delta$
Llama-3 series $\mathcal{L}_{1}$ 0.02 5.94 0.14 -1.57 0.23 1.08 0.29 0.05 5.86 -2.52 -1.68 0.08 -0.97 0.38 38.26 0.87 2.49 26.53 0.37 0.05
$\mathcal{L}_{2}$ 0.64 7.99 0.73 2.45 0.47 -1.08 - 0.00 3.53 0.20 -21.89 0.25 -0.97 - 0.53 0.89 2.19 0.92 0.41 -
$\mathcal{L}_{3}$ 5.93 0.54 - 0.30 -1.06 0.15 - 3.87 0.53 - 0.34 -0.98 -0.05 - 0.80 2.5 - 0.22 0.07
Qwen-2.5 series $\mathcal{L}_{1}$ 0.01 4.32 0.20 -3.73 0.21 -1.17 0.22 -0.58 7.01 -1.89 0.38 0.10 -1.28 0.16 1.85 0.93 0.32 -0.12 0.10 0.17
$\mathcal{L}_{2}$ 0.02 4.78 0.62 4.08 0.32 -1.17 - -0.01 5.84 -0.65 -1.58 0.18 -1.28 - 1.52 0.75 0.92 0.15 0.18 -
$\mathcal{L}_{3}$ 4.77 0.87 - 0.36 -1.15 0.16 - 5.95 -0.91 - 0.16 -1.28 0.02 - 0.76 2.41 - 0.16 0.09
Table 5: Parameter values of fitted parameterizations for P2 Law fitting.
LLM Parameterizations Depth pruning Width pruning 2:4 semi-structured pruning
Llama-3 series $\mathcal{L}_{1}$
$\mathcal{L}_{2}$$\times$ $\times$ $\times$
$\mathcal{L}_{3}$$\times$
Qwen-2.5 series $\mathcal{L}_{1}$
$\mathcal{L}_{2}$$\times$ $\times$ $\times$
$\mathcal{L}_{3}$
Table 6: Compliance of $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ with Condition 2.
(a) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.2-1B pruned by width pruning.
(b) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.2-3B pruned by width pruning.
(c) Loss curves derived by P2 Law and the actual checkpoints of Llama-3.1-8B pruned by width pruning.
Figure 15: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by width pruning.
(a) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-0.5B pruned by depth pruning.
(b) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-1.5B pruned by depth pruning.
(c) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-3B pruned by depth pruning.
Figure 16: Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5 series models pruned by depth pruning.
(a) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-0.5B pruned by width pruning.
(b) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-1.5B pruned by width pruning.
(c) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5-3B pruned by width pruning.
Figure 17: Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5 series models pruned by width pruning.
(a) Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by 2:4 semi-structured pruning.
(b) Loss curves derived by P2 Law and the actual checkpoints of Qwen-2.5 series models pruned by 2:4 semi-structured pruning.
Figure 18: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 seires and Qwen-2.5 series models pruned by 2:4 semi-structured pruning.
(a) Histogram of the normalized importance scores for the embedding channels of Llama-3.2-1B.
(b) Histogram of the normalized importance scores for the embedding channels of Llama-3.2-3B.
(c) Histogram of the normalized importance scores for the embedding channels of Llama-3.1-8B.
Figure 19: Histogram of the normalized importance scores for the embedding channels of Llama-3 series models.
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-1B pruned by depth pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-3B pruned by depth pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.1-8B pruned by depth pruning).
Figure 20: Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by depth pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-1B pruned by width pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-3B pruned by width pruning).
Figure 21: Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by width pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-0.5B pruned by depth pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-1.5B pruned by depth pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by depth pruning).
Figure 22: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by depth pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-0.5B pruned by width pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-1.5B pruned by width pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by width pruning).
Figure 23: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by width pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by 2:4 semi-structured pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by 2:4 semi-structured pruning).
Figure 24: Loss curves derived by P2 Law of models pruned by 2:4 semi-structured pruning for validating generalization on dataset size.
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Llama-3 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by depth pruning).
(c) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by width pruning).
Figure 25: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size.
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models pruned by width pruning).
Figure 26: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models).
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by width pruning).
Figure 27: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models).